SHORTCUT METHOD TO FIND AREA OF TRIANGLE IF ITS THREE SIDES ARE GIVEN


ย 

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Aman Singh


If the side is given

then check if square of any side is equal to the sum of square of other two side .c2=a2+b2

then triangle is right angel triangle

area can be calculate by  1/2 a* b

 if all the side are equal then area =√3/4 *a2


other wise you can use the formula

area=√s(s-a)(s-b)(s-c)

 where a,b,c are side of triangel  and s=(a+b+c)/2


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin